OpenStudy (anonymous):

Solve the equation. q – 29 = –57 –28 28 86 –86<<<< this one

OpenStudy (anonymous):

-28

OpenStudy (anonymous):

how

OpenStudy (espex):

You need to move your -29 to the other side of the equation and do the math to get your answer.

OpenStudy (anonymous):

easy honey q=-57+29 => q=-28

OpenStudy (anonymous):

@eSpeX is -28

OpenStudy (espex):

@heyheyhelpme That is correct, good job.

OpenStudy (anonymous):

Solve the equation. b over 6 = –3 –18 18 1/2 -1/2

OpenStudy (anonymous):

is it 18

OpenStudy (espex):

\[\frac{b}{6}=-3\] When you move the 6 to the other side, what do you get?

OpenStudy (anonymous):

18

OpenStudy (espex):

\[\frac{b}{6}=-3 \rightarrow b=-3*6\] What is -3 times 6?

OpenStudy (anonymous):

-18

OpenStudy (espex):

Exactly, good job.

OpenStudy (anonymous):

@Mertsj what do you think?

OpenStudy (anonymous):

what do you think @.Sam.

OpenStudy (anonymous):

@lopez13

OpenStudy (anonymous):

its correct why you dont believe him

OpenStudy (anonymous):

-18 is the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve the equation. 6r = –54 –60 –48 9 –9

OpenStudy (anonymous):

similar as in previous questions, you have to somehow remove that 6 near r, how you can do that?

OpenStudy (anonymous):

the answer is -9 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

15. Sarah owns a small business. There was a loss of $19 on Thursday and a loss of $12 on Friday. On Saturday there was a loss of $11, and on Sunday there was a profit of $15. Find the total profit or loss for the four days. (1 point)$35 loss $19 profit $27 loss $57 profit

OpenStudy (anonymous):

$27 loss right @Tomas.A

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

15. Sarah owns a small business. There was a loss of $19 on Thursday and a loss of $12 on Friday. On Saturday there was a loss of $11, and on Sunday there was a profit of $15. Find the total profit or loss for the four days. $35 loss $19 profit $27 loss<<<<<<<<this one $57 profit

OpenStudy (anonymous):

what do you thank @zerosniper123

OpenStudy (anonymous):

what do you thank @Tomas.A

OpenStudy (anonymous):

i told you already yes..

OpenStudy (anonymous):

okay bye

OpenStudy (anonymous):

Yes you are correct @heyheyhelpme

OpenStudy (anonymous):

Good job btw! :) Lol